An aqueous solution contains 0.87 g/L of an organic compound. The osmotic pressure of
this solution is 83 mmHg at 29°C. What is the molar mass of the solute?
1
Expert's answer
2020-04-23T13:06:38-0400

P=83mmHg

T=302.15k

C=0.87g/L

"PV=RT"

V=(8.314*302.15)/83=30.26cm3

"M= C*V"

m=0.87*30.26=26.33g